OpenStudy (anonymous):

how can i prove this a/b∙c/d=ac/ad? this is a theorem. can somebody help me on this pls?

OpenStudy (radar):

Unless you've copied incorrectly, it can't be proved, in fact it is false. Just test it with some numbers a=2 b=3 c=5 d=4 a/b x c/d =ac/ad 2/3 x 5/4 = (2x5)/(2x4) 5/6 =5/4 don't think so

OpenStudy (anonymous):

Radar, you made a little mistake. On the denominator RHS, you put 2 instead of 3.

OpenStudy (anonymous):

I need just reasons to prove that why ac/ad is true about that matter?

OpenStudy (anonymous):

You can prove it by inserting in numbers. Radar attempted to that but made a little mistake and came to wrong conclusion.

OpenStudy (anonymous):

Inserting mumbers is a good test, not a proof though...

OpenStudy (anonymous):

example on this is like this: if a=b+c then a-c=b here is the proof: a+(-c)=(b+c)+(-c) the reason on this is ADDITIVE PROPERTY OF EQUALITY a+(-c)=b+{c+(-c)} the reason on this is ASSOCIATIVE OF ADDITION a+(-c)=b+0 the reason on this is INVERSE ADDITION a+(-c)=b the reason on this is IDENTITY OF ADDITION a-c=b the reason on this is SUBTRACTION this is how it goes like !!

OpenStudy (anonymous):

Well, you would need to explain my multiplication, division. Dig up that explanation from your notes.

OpenStudy (radar):

Don't think so, look at the problem as stated at the very top.

OpenStudy (radar):

RHS is ac/ad according to the post ac=10 ad=8 ac/ad=10/8=5/4 which is not 5/6

OpenStudy (radar):

He wants to prove that (a/b) x (c/d) =(ac/ad) which also is c/d (the a cancels)

OpenStudy (radar):

I just can't see it, and your proof is good for what its proving, but it doesn't relate to the problem as stated.

OpenStudy (anonymous):

OK, sorry, Radar, it seems Candice made a mistake in writing down the problem.

OpenStudy (radar):

I was thinking the same thing in fact I mentioned that in the first post. No problem. I realize testing with numbers does not prove, but if it were true, I would come up with something like a=a (I think you know what I mean)